OpenStudy (anonymous):

Find all solutions to the equation 1 = 2cos Θ.

OpenStudy (alekos):

cos(theta) = 1/2 so theta = pi/3 + 2kpi (k=0,1,2....) and theta = -pi/3 + 2kpi (k=0,1,2...)

OpenStudy (anonymous):

I still don't get what I'm suppose to do

OpenStudy (anonymous):

@alekos

OpenStudy (alekos):

Find theta

OpenStudy (princeharryyy):

u need to find theta.. cos(theta) = 1/2 => theta = 60deg and the general solutions are 2npi +or- theta => 1solution is 60 2nd = 2*1*180 + 60 =420deg 3rd = 2*2*180 + 60 = 780deg and so on similarly, for - 2*1*180 - 60 = 300deg 2*2*180 - 60 = 720 and so on.... n =0,1,2,3...

OpenStudy (princeharryyy):

@mashlyashela